### v2

# Synchronized motion of gold nanoparticles in an optothermal trap

Abstract
Optical tweezers have revolutionized particle manipulation at the micro- and nanoscale, playing a critical role in fields such as plasmonics, biophysics, and nanotechnology. While traditional optical trapping methods primarily rely on optical forces to manipulate and organize particles, recent studies suggest that optothermal traps in surfactant solutions can induce unconventional effects such as enhanced trapping stiffness and increased diffusion. Thus, there is a need for further exploration of this system to gain a deeper understanding of the forces involved. This work investigates the behaviour of gold nanoparticles confined in an optothermal trap around a heated anchor particle in a surfactant (CTAC) solution. We observe unexpected radial confinement and synchronized rotational diffusion of particles at micrometre-scale separations from the anchor particle. These dynamics differ from known optical binding and thermophoretic effects, suggesting unexplored forces facilitated by the surfactant environment. This study expands the understanding of optothermal trapping driven by anchor plasmonic particles and introduces new possibilities for nanoparticle assembly, offering insights with potential applications in nanoscale fabrication and materials science.
Unknown University] Department of Physics, Indian Institute of Science Education and Research Pune, Pune, India
1 Introduction
Optical tweezers have revolutionized our ability to control and study particles at the microscale and nanoscale, making them invaluable in fields like biophysics, nanotechnology, and material science. 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11 This technique uses focused laser beams to trap, move, and manipulate microscopic particles, such as biological cells, colloidal microparticles and nanoparticles. Optical trapping can easily adapt and integrate with various materials and geometries to improve control and efficiency. For instance, plasmonic trapping harnesses the resonance of conduction electrons in metallic nanoparticles to create stronger localized fields, enabling more efficient traps and enhanced precision at smaller scales.12, 13, 14, 15, 16, 17, 18, 19 Optical trapping near a surface, like a glass wall, can also significantly alter trap efficiency and dynamics.20, 21, 22 These surface-mediated effects are critical for understanding particle-surface interactions, with applications like optical nanomotors23, 24, 25, 26, 27, 28 and optical printing.29, 30, 31 Optical tweezers can also induce optical binding between particles. This light-scattering mediated interaction can be used to manipulate multiple particles into stable self-assembled structures such as chains or arrays.32, 33, 34, 35, 36, 37 Conventional optical binding uses a large illumination area, but recent studies show indirect optical binding outside the focal spot of the laser. 38, 39, 40
Optical tweezers can also utilize the localized heating of plasmonic particles and surfaces. 41, 42, 43, 44, 45, 46, 47 A thermal gradient at the microscale can help facilitate thermophoresis, thermo-osmosis, thermodiffusion, and thermoelectricity, among other auxiliary fields. These auxiliary fields provide additional control for particle manipulation at nanoscale. Optothermal traps enable particle confinement at much lower laser powers than traditional optical tweezers, reducing the risk of damage to heat sensitive materials and studying the interaction, dynamics, and assembly of different kinds of particles.48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62 In surfactant solutions, optothermal trapping has achieved even more efficient trapping at low powers. This is also referred to as opto-thermoelectric (OTE) trapping. By leveraging optothermal manipulation and surfactant-mediated interactions, it is possible to create novel nanoparticle assemblies at lower powers with varied applications.63, 64, 65, 30
Optothermal trapping has enabled diverse and efficient particle manipulation at the nanoscale. However, the mechanisms governing particle behaviour in surfactant-enhanced optothermal traps are not completely understood. While optothermal forces can attract and organize particles, introducing surfactants fundamentally changes trapping dynamics, creating unique interactions and behaviours not observed in standard optothermal or optical binding setups. In this study, we report an unconventional trapping behaviour in such an environment.
We show the trapping of gold nanoparticles in an anchor particle-assisted trap in a surfactant solution. The observed dynamics are novel as particles get trapped outside the beam focus, and multiple trapped particles show synchronized diffusive motion with micrometre separation between them. We have previously demonstrated the ability of anchor particle-assisted traps to efficiently trap individual nanoparticles and nanoparticle assemblies.54, 66 In this study, we focus on the dynamics of particles in anchor particle optothermal trap. We find that particles in solution are trapped at a radial position between 1 and 2 microns away from the drop-casted particle. Furthermore, trapped particles undergo rotational diffusion around the drop-casted particle. This rotational diffusion is highly synchronised for various particles, even with separations of an order of microns between the particles. As discussed in the Results section, this synchronisation is neither due to temperature-induced thermophoresis nor optical binding. We conducted numerical simulations to investigate the influence of various inter-particle forces. Our experimental findings also demonstrate that trapped particles can spontaneously exhibit directed rotations.
Figure 1: Schematic figure. a) The optical schematic shows the dual-channel optical microscope setup, which consists of two objectives (100$\times$, 0.95 NA and 100$\times$, 1.4 NA) for excitation and signal collection, respectively. b) Schematic figure of the optothermal trap using 400 nm gold nanoparticles (AuNP) in the presence of surfactant (CTAC). c) Transmission Electron microscopy images of the used AuNPs. d) Experimental data corresponding to Supplementary Video 1 showing the trapping away from the centre and synchronised diffusion along the circumference. e) Angular trajectories of the trapped NPs showing synchronised rotational diffusion.
2 Results and discussion
We investigate the optothermal trapping of gold nanoparticles in surfactant solution on a heated anchor particle-based optothermal trapping platform. To resolve the dynamics compared to our previous studies on trapping in surfactant solutions, we have performed the experiments on a dual channel setup as shown in the optical schematic in Figure 1a. The setup is configured to collect light from the sample plane using a 100$\times$, 1.49 NA objective lens and captures images with a fast camera at 500 frames per second. The schematic Figure 1b shows the experiment geometry. As described in the methods section, a 400 nm gold particle is drop-casted on a glass substrate and is firmly attached to the cover glass. We illuminate this particle, called the anchor particle, with the 532 nm laser using a high numerical aperture objective lens (100$\times$, 0.95). The 400 nm AuNP has a significant absorption cross section at the used laser wavelength and leads to heating. The heated particle sets a temperature gradient on the glass substrate and the surrounding fluid, which can, in turn, set up a thermo-osmotic slip flow in the surrounding fluid. 67, 52, 49, 68 Gold nanoparticles diffusing in a surfactant solution follow the fluid flow and come near the anchor particle and get confined at a radial distance as mentioned in the introduction and shown in figure 1d.
Figure 1d shows the transmission electron microscopy images of the 400 nm gold nanoparticles used for experiments. Figure 1d shows representative experimental observations corresponding to our schematic. The AuNP in the centre is drop-casted, and the AuNPs in the solution get trapped at around a micrometre distance far from the centre of the beam. In Figure 1d, the two trapped particles perform rotational diffusion at a constant radial separation from the anchor particle. Figure 1e shows rotational diffusion in angular coordinates around the particle. It can be observed that the diffusion is synchronized as the angular trajectories move with each other. We now proceed to describe the role of surfactant and anchor particle in the optothermal trapping and discuss the observed synchronization. We will also discuss the role of forces involved in the trapping process in the following subsections.
2.1 Anchor particle driven trapping
Figure 2: Dynamics of gold nanoparticles in different traps (corresponding to the Supplementary video 2): a) without anchor gold particle, c) without surfactant (CTAC), and e) with both anchor particle and surfactant. The tracks of particles are shown for each condition in b), d), and f), respectively.
Recall that the essential elements of the trap we observe are the anchor particle and the surfactant. We attempt to decompose the trapping process by considering the effects of these essential elements of the trap individually. We discuss these results systematically in figure 2. First, we discuss the trapping behaviour of particles in the surfactant solution without the anchor particle. The surfactant used in the experiments was Hexadecyltrimethylammonium chloride (CTAC), and the concentration used was always five millimolars (mM). The images are taken in transmission brightfield configuration as described in methods. In figure 2a, the gold particles in the surfactant solution are shown confined around the laser beam. In this case, the beam is made partially visible and creates a dark spot in the centre of the images in figure 2a. The trajectories of the particles are random and Brownian, as shown in figure 2b, with some attraction towards the beam. The particles hover near the beam but do not enter the high-intensity region of the beam owing to the transverse optical scattering force.
Secondly, Figure 2c shows the trap with the anchor particle without the surfactant. Here, the anchor particle heats up and sets up fluid flow that drives the particles in out-of-plane circuitous trajectories. The particles come radially inwards towards the anchor particle, diffuse out of the plane, and move radially outwards with the fluid flow. The particles again come in the sample plane at a larger radial distance and repeat the cycle. The trajectories of all particles correspond to this behaviour in figure 2d.
Finally, in figure 2e, we show the behaviour of trapped particles around the anchor particle in the presence of 5 mM CTAC. The AuNPs show remarkably different behaviour as different particles show confinement at various radii. The particles get confined radially and diffuse rotationally around the circumference. This is visible prominently in trajectories in figure 2f. In all three cases, the particles are not trapped in the centre of the intensity gradient or the heat gradient, as observed in previous studies. We believe this is because of the strong optical scattering force on the 400 nm AuNPs from the 532 nm laser beam. The confinement behaviour of diffusing AuNPs at some fixed radii from the anchor AuNP is only observed in the presence of surfactant in the solution. The behaviour is absent if we remove either the surfactant or the anchor particle, as is shown in figure 2a-d. A comparison of the dynamics of particles around the anchor particle with and without surfactant is presented in Supporting Information 1.
2.2 Synchronised motion of trapped particles
Figure 3: Synchronised motion of trapped gold nanoparticles. a) Angular positions of trapped particles show synchronous displacements, as is also visible in Supplementary video 2. b) shows the distribution of linear separation d12, d23, and d34 between the particles P1, P2, P3 and P4, respectively, as shown in the inset. The top four trajectories in a) correspond to P4, P3, P2, and P1.
In the previous subsection, we described the offset radial confinement and azimuthal diffusion of a gold nanoparticle around the anchor particle. When multiple particles are present in this trap, they show rotational diffusion around the circumference, which is synchronised, as shown in Figure 3a. It shows the angular positions of the trapped particles in Figures 2e and f for around 12 seconds. The azimuthal diffusion is not completely coupled but varies slightly as the angular position tracks sometimes come closer. Figure 3b shows the distribution of consecutive linear distances d12, d23, and d34 between particles P1, P2, P3 and P4, as marked in the inset. The particle P1 is the last to join the assembly, and distances are measured after it has settled at a radial position. The histograms of the distances show that the linear distance between particles is not fixed but has peaks. This suggests that there is a tendency to maintain a fixed separation from the surrounding particles, but there are also some deviations from that tendency, presumably due to thermal noise and local surface conditions. The consecutive interparticle distances range between 0.8 to 1.6 microns in this case.
This synchronised diffusion of particles with a micrometre order separation suggests some repulsive force of the micrometre range between particles. Other reports have shown such repulsion in different colloids due to the thermophoretic repulsion between heated particles51 or the optical binding force35. We suspect neither of these is responsible for the repulsion in this case. Thermophoresis is the tendency of particles to migrate in a temperature gradient.69, 70, 71 It is a surface phenomena and the existence of a temperature gradient on the surface of the particle is important to observe thermophoresis.72 But owing to the high thermal conductivity of the gold nanoparticles, they have a uniform temperature across their surface. Consequently, they do not show thermophoretic repulsion. Moreover, thermophoretic repulsion creates assemblies where particles are equispaced around the circumference, as shown in a previous report51. In the current study, the particle separation is not uniform, as shown in figure 3b. See Supplementary Video 3 and Supporting Information 2 for further discussion. Meanwhile, for repulsion due to optical binding, which is a light scattering mediated interaction, the distance between the particles is integer multiples of the half-wavelength of the laser beam.33 It is shown in figure 3b that the distance does not conform to that condition. Furthermore, our trapping beam has relatively low power (< 15 mW) and is focused on the anchor particles. The trapped particles are excluded from the high-intensity beam spot region, which has a diameter of 1 micron. This contrasts with the broad beam used with higher powers for optical binding experiments, as shown in the literature.
2.3 Numerical estimation of forces
Figure 4: Computation of Forces. a) Optical potential around the anchor AuNP calculated by integrating the optical scattering force. b) Temperature distribution around the heated anchor nanoparticle. c) The computed Optical, Opto-thermoelectric, and Thermo-osmotic slip induced drag forces on a particle near the anchor particle at various distances. The sum of all these forces along the x direction is also plotted to show the attractive behaviour at large distance.
In this section, we study numerically various forces present in our system which could give rise to the observed dynamics of particles. We focus on the observation of the confinement at a radial position. The major forces present in the system are the optical forces due to the laser beam, heating-induced opto-thermoelectric force, and thermo-osmotic slip flow-induced drag force. In the presence of the laser beam, the diffusing nanoparticles experience a force due to momentum exchange with the optical field. The resultant optical force is generally decomposed into scattering and gradient force, where the gradient force is responsible for the stable trapping of particles at the beam centre. In our study, the AuNPs used in the solution have a large scattering cross section and experience a large scattering force. Consequently, they are unable to access the high-intensity region of the beam. We show the optical potential profile around the anchor particle, which is shown in Figure 4a. The potential is calculated at 1 mW of illumination power with a beam waist of around 1 micron. The optical potential is positive at all places, indicating that the in-plane force on any diffusing particle is always repulsive, pointing away from the anchor particle. The optical potential is calculated using the generalized multiparticle mie theory-based calculation as described in the methods. The potential profile is not circularly symmetric as the excitation laser is x-polarised. This gives a large force along the y-direction as compared to the x-direction.
We also simulate the heating of the anchor particle by calculating the absorption coefficient of the gold nanoparticle using the Mie theory as shown in Figure 4b. Our simulated temperature rise matches well with the experimental measurement (See Supporting Information 3) of the heating of anchor nanoparticles. We can then calculate the heating of the surrounding substrate and fluid, allowing us, in turn, to calculate the fluid flow around the particle. This fluid flow is classified into two parts. The first part is convection, caused by changes in the density of the heated fluid. The second part is thermo-osmotic slip flow, resulting from differences in the free energies of the fluid across a temperature gradient parallel to the substrate. Since thermo-osmotic slip flow is the dominant contribution, we will focus on that. We modelled this thermo-osmotic slip flow using a commercial FEM-based solver using the expression from the literature $v_{||}=\chi\frac{\nabla_{||}T}{T}$, where $\chi$ is the thermo-osmotic coefficient, $T$ the temperature and $\nabla_{||}T$ is the temperature gradient parallel to the surface. We then calculate the drag force on the particle from this fluid flow using the stokes drag formula $F_{slip}=6\pi\eta Rv_{||}$, where $\eta$ is the viscosity of the fluid and $R$ is the radius of the particle.
We can also calculate the opto-thermoelectric (OTE) force using the temperature profile around the particle. OTE force has been shown to be very prominent in optothermal trapping studies in the presence of surfactant. The OTE force arises in systems with surfactants and local heating, as shown in the literature. The working principle of this force is the separation of charges of the cationic surfactant in the solution. The surfactant (CTAC) molecule dissociates into the chloride ion and the CTA+ cation in the solution. For concentrations exceeding the critical micellar concentration of the surfactant, the CTA+ ions self-assemble into micelles. The CTA+ ions are also adsorbed on all particle surfaces, rendering them positively charged. In the presence of the temperature distribution, the macro-cations (micelles) and the anions start moving down the temperature gradient owing to their intrinsic thermophoresis. However, the drift speed of migration of the macro-cations is larger than the speed of anions as the anions are smaller and more diffusive. This difference in the thermo-diffusivities results in a separation of charges in the steady state. This separation of charges creates an electric field which drives the positively charged particles towards the heat centre. This mechanism can provide a large force for stable trapping in many cases, but it is not significant in our case, as can be seen in Figure 4c. The OTE force is directly proportional to the product of the temperature gradient and temperature. The temperature field decays as $\frac{1}{r}$ and so the temperature gradient decays as $\frac{1}{r^{2}}$ and thus the opto-thermoelectric force scales as $\frac{1}{r^{3}}$. Thus, the OTE force reduces in magnitude significantly compared to the other forces in the separation range of interest.
We show a combination of all computed forces in Figure 4c, which shows an equilibrium position near 0.6 microns. We note that since the contribution of the OTE force is negligible, the force profile can be almost identical for the case without the surfactant, which has starkly different dynamics. We conclude this section by mentioning that while simulations show the force profile to explain the observations, they are incomplete in explaining the repulsion and synchronised motion between particles.
2.4 Emergence of spontaneous rotation
Figure 5: Spontaneous rotation of trapped particles. a) Time series of particles showing the emergence of spontaneous rotation in trapped particles. b) Trajectories of particles.
In this section, we experimentally demonstrate that spontaneous rotation can emerge in the trapped particle assemblies. Figure 5a shows the time series of the particles corresponding to the supplementary video 4. The particles are trapped at radial positions, as described previously. The three-particle system, as shown in figure 5a, diffuses and gets into a particular configuration due to thermal noise. After achieving a particular configuration, the particle system starts rotating suddenly. The particle system rotates for two rotations in about 5 seconds. Then, the particles rearrange again due to fluctuations, and the rotation ceases. The position distribution in Figure 5b shows that the particles cover a circular trajectory.
3 Methods
3.1 Materials
Milli Q water was used to prepare the sample. Cetyltrimethylammonium chloride (CTAC) was obtained from Sigma-Aldrich in solution (25 wtin water). Gold nanoparticles with a 400 nm diameter, stabilised suspension in citrate buffer, were purchased from Sigma-Aldrich. We have characterised the purchased particles through TEM images and UV-Vis spectra, which are shown in supplementary information. Two kinds of suspensions were created with gold nanoparticles. One in water and the other in ethanol. 1 mL of nanoparticle suspension was centrifuged, and the supernatant was removed and redispersed in 2 mL of ethanol for drop-casting on the surface. Similarly, 1 mL of nanoparticle suspension was centrifuged and redispersed in 1 mL of MilliQ water. Two 50 µL solutions were prepared in micro-centrifuge tubes using MilliQ water to prepare nanoparticle suspension in surfactant solution. One with 5 µL of previously prepared aqueous gold NP suspension added to 45 µL water. Another tube had 10 µL of freshly prepared 50 mM CTAC concentration solution added to 40 µL of water. These 50 µL solutions were mixed to give the final solution, which was used for trapping. Goldseal Cover Glass from Ted Pella, Inc. were used for experiments after cleaning with acetone. A double-sided adhesive spacer from Grace Bio-Laboratories (SecureSeal imaging spacer, 120 µm thickness) was used to build a microfluidic chamber as follows. The spacer is pasted on the cover glass. 2 µL of the alcoholic gold suspension was drop-casted on the glass coverslip and allowed to dry by evaporation for 20 minutes. This gives gold NPs which are firmly attached to the surface. After 20 minutes, a 10 µL solution of the surfactant mixed with aqueous suspension is poured into the spacer, and the chamber is sealed with another cover glass from the top.
3.2 Optical Measurements
The experiments were performed on a custom-built dual-channel microscope. The imaging was performed in brightfield mode, and trapping was performed in upright geometry. A 100$\times$, 0.95 NA objective lens (Olympus MPlanApo N) focused the laser from the top of the AuNP anchored on the cover glass. The same objective also illuminates the sample plane with white light from the top. Another 100$\times$, 1.49 NA objective lens (Olympus UApo N, oil immersion) was used to collect all the transmitted light and the images were captured using a fast camera with a frame rate of 500. The laser wavelength is filtered from the collected light before reaching the camera using a notch filter and a long pass filter. The setup was equipped with a three-dimensional piezoelectric translation stage for high-precision control of the sample.
A 532 nm linearly polarised Gaussian laser beam was used for the experiments. The power of the laser was controlled using a half-waveplate and a polarising beam splitter. This assembly provides the capability to control the power continuously and measure the power in the sample plane by calibration. The laser powers used are less than 15 mW in the sample plane. The beam was not expanded, and the back aperture of the objective was underfilled.
3.3 Particle Tracking
A fast camera recorded the particle trajectories at a frame rate of 500 frames per second. The images were tracked using the Trackmate plugin of the FiJi distribution of ImageJ73, 74. The images shown were taken using brightfield illumination, and the gold nanoparticles are shown as dark objects on a bright background. These images were then inverted for tracking using TrackMate plugin. The images extracted from the tracked processed video show particles as white objects on a black background.
3.4 Numerical Calculations
We calculated the scattering forces between particles using a Python-based generalized multiparticle mie theory package, MiePy. The forces were then integrated to get the potential profile (see Supporting Information 5). The absorption cross-section of the nanoparticle was calculated using Mie theory, which was used to calculate the temperature rise of the nanoparticle. The temperature distribution around the nanoparticle was then calculated by solving the Heat-diffusion equation by convolving the heat source density with the Green’s function. Using the temperature rise of the nanoparticle, we conducted numerical simulations using a finite element method (FEM) solver to calculate the thermo-osmotic slip flow-induced force (see Supporting Information 6). The OTE force is similarly computed using the temperature profile as shown in Supporting Information 7.
4 Conclusion
This study investigates the dynamics of nanoparticles in anchor particle-based optothermal traps, focusing on the behaviour of gold nanoparticles in surfactant solutions. We observed that particles are trapped some distance from the centre and exhibit synchronized rotational diffusion and spontaneous rotation around a heated anchor particle. The dynamics suggest the presence of some repulsive interparticle force, which helps synchronise the particles’ motion even with interparticle separation of around 1 micron. These unique dynamics open possibilities for precisely tuning particle interactions, which could be highly valuable for applications in nanotechnology, such as assembling nanoparticle arrays or manipulating particles in nanoscale systems.
The presence of surfactants, particularly in the form of CTAC solutions, plays a critical role in influencing the trapping and motion of nanoparticles. Our results suggest that surfactants help confine particles to specific radii around the anchor particle, leading to organized rotational diffusion. This behaviour was absent when either the anchor particle or the surfactant was removed, underscoring its importance in achieving controlled trapping. Overall, our study aims to contribute to a deeper understanding of particle dynamics in complex trapping environments, offering insights that could be used to design new nanoparticle assembly techniques. These findings hold potential for advancing applications in materials science, nanofabrication, and optoelectronics, where controlled particle interaction and organization are important. {acknowledgement} The authors thank Dr Adarsh Vasista, Dr Sunny Tiwari and Dr Diptabrata Paul for their valuable discussions on this project. We acknowledge using common experimental facilities in the Chemistry Department at IISER Pune. AS acknowledges the Ministry of Education, Government of India, for the Prime Minister Research Fellowship. This work was partially funded by AOARD (grant number FA2386-23-1-4054) and the Swarnajayanti fellowship grant (DST/SJF/PSA-02/2017-18) to G.V.P.K.
{suppinfo}
Supplementary information containing the following information and videos can be found at this link.
Comparison of particle distribution and dynamics with and without CTAC. Two particle synchronized rotational diffusion. Measurement of temperature using liquid crystal phase transition. Computation of temperature rise and distribution. Computation of optical potential. Computation of thermo-osmotic slip flow and force. Computation of opto-thermoelectric force. (pdf)
Supplementary Video 1: Experimental video showing trapping and synchronization of gold particles. Supplementary Video 2: Comparision of trapping dynamics in various conditions. a) without anchor particle and with CTAC, b) without CTAC and with anchor particle, c) with anchor particle and 5 mM CTAC. Supplementary Video 3: Experimental video with simultaneous angular position graph showing synchronised motion. Supplementary Video 4: Experiment video showing spontaneous rotation in trapped particle assembly.